# bHUN-REN Wigner Research Centre for Physics, Konkoly-Thege Miklos ut 29-33, 1121 Budapest, Hungary cInstitut de Physique Théorique [^1] Unité Mixte de Recherche 3681 du CNRS, Université Paris Saclay, CNRS, 91191 Gif-sur-Yvette, France

### v2

@@ 1 Introduction and summary
Supersymmetric gauge theories with extended supersymmetry provide a powerful framework for studying nonperturbative aspects of quantum field theory. In three dimensions, a prominent example is the ABJM model Aharony:2008gk; Aharony:2008ug. It describes two supersymmetric Chern–Simons theories with gauge group $U(N)$ and opposite levels, $k$ and k-k, coupled to four matter supermultiplets in the bifundamental representation of $U(N)\times U(N)$. The theory exhibits $\mathcal{N}=6$ superconformal symmetry which, for the special values $k=1$ and $k=2$, is enhanced to $\mathcal{N}=8$, the maximal amount of supersymmetry allowed in three dimensions.
@@
At large $N$, two qualitatively different regimes emerge, depending on the hierarchy between $N$ and $k$. When both parameters are large, observables can be systematically expanded in powers of $1/N^{2}$ by keeping the effective ’t Hooft coupling $\lambda=N/k$ fixed. Within the AdS/CFT correspondence Aharony:1999ti, this limit corresponds to the perturbative genus expansion of a type IIA superstring propagating on $\mathrm{AdS}_{4}\times\mathbb{CP}^{3}$ Benna:2008zy; Arutyunov:2008if; Stefanski:2008ik; Hosomichi:2008jb.
@@
Supersymmetric localization Pestun:2016zxk provides a powerful framework for studying different regimes of the ABJM theory. For a special class of observables preserving supersymmetry, this technique allows us to reduce the infinite-dimensional path integral to a finite-dimensional matrix integral. Originally developed for four-dimensional supersymmetric Yang-Mills theories Pestun:2007rz, it was later generalised to three-dimensional supersymmetric Chern–Simons theories Kapustin:2009kz.
@@
In the ABJM theory, the matrix model obtained via supersymmetric localization provides a nonperturbative representation of a wide class of observables, including the partition function on the three-sphere $S^{3}$, circular Wilson loops Drukker:2008zx; Chen:2008bp; Drukker:2009hy; Drukker:2019bev and the bremsstrahlung function Lewkowycz:2013laa; Bianchi:2014laa; Bianchi:2017ozk; Bianchi:2018bke; Bianchi:2018scb. Solving this matrix model and deriving exact expressions for the observables that are valid for arbitrary values of parameters of the ABJM theory is an extremely difficult and challenging task.
@@
A powerful technique for addressing this problem is the Fermi gas approach Marino:2011eh. In this framework, the ABJM matrix model is reinterpreted as a one-dimensional quantum gas of interacting fermions, with the Chern–Simons level $k$ playing the role of an effective Planck constant, $\hbar=2\pi k$. This approach has been successfully applied to the computation of the partition function $Z(N,k)$ of the ABJM theory on the round three-sphere $S^{3}$, see Hatsuda:2015gca; Marino:2016new for reviews. Remarkably, the generating function $\Xi(z,k)=\sum_{N\geq 0}z^{N}Z(N,k)$, which has the meaning of the grand canonical partition function of the Fermi gas, admits a compact representation as a Fredholm determinant of a certain integral operator.  [^2] It is noteworthy that various observables in four-dimensional maximally supersymmetric Yang-Mills theory can likewise be expressed in terms of a Fredholm determinant of an integrable Bessel kernel (see Beccaria:2022ypy and references therein). This representation is particularly powerful, as it enables the computation of these observables for arbitrary values of the ’t Hooft coupling Bajnok:2024epf; Bajnok:2024ymr; Bajnok:2024bqr. In the strong-coupling regime, corresponding to the limit $\hbar=2\pi k\to 0$, the Fredholm determinant can be evaluated using a semiclassical expansion. When supplemented with nonperturbative (instanton) corrections, this analysis leads to the expression for the partition function in terms of Airy function.
@@
The Fermi gas formalism can also be applied to the computation of supersymmetric Wilson loops. In the ABJM theory, one can define two distinct supersymmetric Wilson loops, $W^{1/6}$ and $W^{1/2}$, which preserve four and twelve superconformal symmetries, respectively, out of the total $24$ of the theory Drukker:2008zx; Chen:2008bp; Drukker:2009hy; Drukker:2019bev. The $1/6$ BPS Wilson loop is defined as the vacuum expectation value
@@
Both Wilson loops (1) and (2) depend on a nonnegative integer $n$, which specifies the winding number. For $n=0$, they satisfy the normalization condition $W_{0}^{1/6}=1$ and similarly $W_{0}^{1/2}=1-(-1)^{n}$. As a consequence of (2), it is sufficient to focus on the computation of the $1/6$ BPS Wilson loop (1).
The semiclassical expansion of the Wilson loops (1) and (2) within the Fermi gas approach was derived in Klemm:2012ii. Neglecting nonperturbative (instanton) corrections, the resulting expressions for $W_{n}^{1/6}$ and $W_{n}^{1/2}$ take the form of ratios of Airy functions and their derivatives. In parallel development, the $1/6$ BPS Wilson loop was computed numerically with high precision for a wide range of parameters $k$ and $N$ in Okuyama:2016deu. These numerical studies revealed that, for winding number $n\geq 2$, the perturbative part of the $1/6$ Wilson loop differs from the semiclassical expression obtained in Klemm:2012ii. This discrepancy was unexpected, especially given that the semiclassical expression for the partition function $Z(N,k)$ agrees perfectly with numerical data Hatsuda:2012hm. The origin of the mismatch remained unexplained.
@@
In the Fermi gas framework, the Wilson loop (1) admits an exact representation in terms of the resolvent of the same operator that appears in the Fredholm determinant expression for the partition function $Z(N,k)$. The corresponding Fredholm determinant is known to satisfy a system of coupled nonlinear TBA-like equations. These equations were originally introduced independently in the study of $\mathcal{N}=2$ models in two dimensions Cecotti:1992qh and of self-avoiding 2d polymers Zamolodchikov:1994uw, and were subsequently proven in Tracy:1995ax.
Building on the method developed by Tracy and Widom in Tracy:1995ax, we derive a set of exact equations governing the function $W_{n}^{1/6}$. We show that these equations are sufficiently powerful to determine the Wilson loops for arbitrary values of the parameters, without relying on the semiclassical approximation.
@@
We verified that the obtained results for the Wilson loops are in perfect agreement with the expressions obtained in Okuyama:2016deu from numerical fitting. Combined with the explicit expressions for the free energy, they provide an efficient method for computing the Wilson loops for arbitrary values of the parameters of the ABJM theory.
@@ 2 Fermi gas approach
Supersymmetric localization allows us to express the partition function of the ABJM theory on the three-sphere and the $1/6$ BPS Wilson loop (1) in terms of the same function $W_{n}=W_{n}(N,k)$. This function is given by a finite-dimensional integral Kapustin:2009kz
@@
To illustrate this, it is sufficient to consider the Abelian case with $N=1$. In this case, the integral in (5) can be evaluated explicitly, yielding Okuyama:2016deu
@@
This expression is regular for $k$ obeying (7), but it develops a singularity at $k=2n$. The singularity originates from the divergence of the integral (5) at this value of $k$. As we demonstrate below, the same singular behaviour persists for arbitrary $N$.
@@ 2.1 Generating functions
As follows from the first relation in (6) these two functions are not indepedent and are related to each other as
@@
In the similar manner, the generating function $\mathcal{W}_{n}(z,k)$ has the meaning of the expectation value of the holonomy $\mathop{\rm tr}\nolimits(U^{n})=\sum_{i=1}^{N}e^{n\mu_{i}}$ in a grand canonical ensemble.
@@
Applying (6) and replacing the functions $Z(N,k)$ and $W_{n}(N,k)$ with their integral representation (5), we obtain rather complicated expressions for the generating functions (2.1). As was shown in Marino:2011eh, these expressions can be cast into a compact form by noticing that the function $Z(N,k)$ admits a determinant representation
@@
In a similar way, the generating function $\mathcal{W}_{n}(z,k)$ defined in (2.1) can be written in the operator form Klemm:2012ii; Hatsuda:2013yua,
@@
where the operator $\bm{U}=e^{(\bm{x}+\bm{p})/k}$ provides a representation of the holonomy in the grand canonical ensemble. We verify that for $n=0$ the relations (15) and (17) are in agreement with (10). For $n\neq 0$, the operator in (17) can be simplified using the Baker-Campbell-Hausdorf formula as $\bm{U}^{n}=e^{n\bm{p}\over k}e^{n(\bm{x}+i\pi n)\over k}$ leading to the following representation Okuyama:2016deu
@@
As we demonstrate below, relations (15) and (18) provides a powerful tool for evaluating un-normalized Wilson loop $W_{n}(N,k)$ for arbitrary positive $N$ and $k$. For the first few values of $N$, this function can be found from small $z$ expansion of the generating functions (2.1). Expanding the right-hand side of (15) and (18) in powers of $z$, we get from (2.1)
@@ 2.2 Semiclassical approximation
The advantage of considering this regime is that, within the Fermi gas framework, the small-$k$ behaviour of the generating functions $\Xi(z,k)$ and $\mathcal{W}_{n}(z,k)$ defined in (2.1) can be systematically derived from the semiclassical WKB expansion of the quantum Fermi gas. It is important to emphasize, however, that it is not obvious a priori whether this expansion remains valid for finite values of $k$, or, equivalently, whether the semiclassical expressions can be analytically continued from small $k$ to positive integer values without encountering singularities. This question has been investigated in Hatsuda:2012hm; Okuyama:2016deu by comparing semiclassical predictions with numerical evaluations of (15) and (18). The analysis shows that, while the semiclassical expansion accurately describes the function $\Xi(z,k)$, it fails for $\mathcal{W}_{n}(z,k)$ when $n\geq 1$.
@@
By combining the semiclassical expansion with numerical analysis and imposing certain nontrivial consistency conditions, the grand canonical partition function $\Xi(z,k)$ defined in (2.1) and (15) was conjectured to have the following form at large $z=e^{\mu}$ and finite $k$, see Hatsuda:2015gca for a review,
@@
where the constant term $A(k)$ admits an explicit integral representation Marino:2011eh; Hanada:2012si.
@@
Furthermore, the grand canonical partition function (21) simplifies significantly for $k=1,2,4,8$ and the corresponding function $\Xi(N,k)$ can be written in a closed form in terms of the Jacobi theta functions Codesido:2014oua; Grassi:2014uua; Okuyama:2016xke.
@@
Previous studies have shown that, at large $\mu$ and finite $k$, the function $\mathsf{W}_{n}(\mu,k)$ admits the following general representation Klemm:2012ii; Hatsuda:2013yua; Okuyama:2016deu
@@
The perturbative contribution (29) has been computed using two complementary approaches: through the semiclassical expansion of the Fermi gas ensemble Klemm:2012ii and via high-precision numerical evaluation of the winding Wilson loop Okuyama:2016deu. The corresponding constant term was found to be different
@@
The imaginary and real parts of the nonperturbative contribution to the $1/6$ BPS Wilson loop (30) were studied numerically in Hatsuda:2013yua; Okuyama:2016deu for the first few winding numbers, n=1,2,3,4n=1,2,3,4444In virtue of (2), the function $\mathop{\rm Im}\nolimits\mathsf{W}_{n}^{\text{np}}(\mu,k)$ coincides with the nonperturbative part of the $1/2$ BPS Wilson loop. Because of the normalization factor in (28), the function $\mathop{\rm Re}\nolimits\mathsf{W}_{n}^{\text{np}}(\mu,k)$ corresponds to the imaginary part of the $1/6$ BPS Wilson loop in the conventions of Okuyama:2016deu. By evaluating $\mathsf{W}_{n}^{\text{np}}(\mu,k)$ numerically for some positive integer values of $k$, the exact values of the instanton coefficients $w_{m,\ell}(\mu)$ were guessed for the lowest values of $m$ and $\ell$ through numerical fitting. Building on this analysis, a number of conjectures were formulated concerning the properties of the nonperturbative function $\mathsf{W}_{n}^{\text{np}}(\mu,k)$.
@@ 3 Tracy-Widom approach
This approach was originally developed by Tracy and Widom Tracy:1995ax, who proved the conjecture formulated in Cecotti:1992qh; Zamolodchikov:1994uw, according to which the Fredholm determinant of an integral operator of the same type as in (15) satisfies a TBA-like system of nonlinear equations. The integral operator in (15) involves the kernel $\rho(x,y)$ defined in (14). It represents a special case of the more general class of kernels analyzed in Tracy:1995ax [^5] Note that the kernel $\rho(x,y)$ is not symmetric, $\rho(x,y)\neq\rho(y,x)$. To bring it to the form considered in Tracy:1995ax, it has to be symmetrized by a similarity transformation, which leaves the Fredholm determinant (15) invariant.
We begin by introducing an auxiliary function which plays a central role in our analysis  [^6] In the notation of Tracy:1995ax, this function is given by the sum of functions $Q(\theta)$ and $P(\theta)$ for $\theta=x/k$.
@@
The rational for introducing the auxiliary function (32) is that the generating function (18) can be expressed in terms of this function
@@
Replacing the $\psi$-function in (3) with its small $z$ expansion (3), we can systematticaly expand the generating function $\mathcal{W}_{n}(z,k)$ in powers of $z$ and reproduce the relation (19). To find $\mathcal{W}_{n}(z,k)$ for arbitrary $z$, we have to determine the function (32). This will be done in two steps. First, we formulate a set of nontrivial relations the function $\psi(x|z)$ has to satisfy and, then, present the solution.

### v3

@@ 1 Introduction and summary
Supersymmetric gauge theories with extended supersymmetry provide a powerful framework for studying nonperturbative aspects of quantum field theory. In three dimensions, a prominent example is the ABJM model Aharony et al. (2008b, a). It describes two supersymmetric Chern–Simons theories with gauge group $U(N)$ and opposite levels, $k$ and k-k, coupled to four matter supermultiplets in the bifundamental representation of $U(N)\times U(N)$. The theory exhibits $\mathcal{N}=6$ superconformal symmetry which, for the special values $k=1$ and $k=2$, is enhanced to $\mathcal{N}=8$, the maximal amount of supersymmetry allowed in three dimensions.
@@
At large $N$, two qualitatively different regimes emerge, depending on the hierarchy between $N$ and $k$. When both parameters are large, observables can be systematically expanded in powers of $1/N^{2}$ by keeping the effective ’t Hooft coupling $\lambda=N/k$ fixed. Within the AdS/CFT correspondence Aharony et al. (2000), this limit corresponds to the perturbative genus expansion of a type IIA superstring propagating on $\mathrm{AdS}_{4}\times\mathbb{CP}^{3}$ Benna et al. (2008); Arutyunov and Frolov (2008); Stefanski (2009); Hosomichi et al. (2008).
@@
Supersymmetric localization Pestun and others (2017) provides a powerful framework for studying different regimes of the ABJM theory. For a special class of observables preserving supersymmetry, this technique allows us to reduce the infinite-dimensional path integral to a finite-dimensional matrix integral. Originally developed for four-dimensional supersymmetric Yang-Mills theories Pestun (2012), it was later generalised to three-dimensional supersymmetric Chern–Simons theories Kapustin et al. (2010).
@@
In the ABJM theory, the matrix model obtained via supersymmetric localization provides a nonperturbative representation of a wide class of observables, including the partition function on the three-sphere $S^{3}$, circular Wilson loops Drukker et al. (2008); Chen and Wu (2010); Drukker and Trancanelli (2010); Drukker and others (2020) and the bremsstrahlung function Lewkowycz and Maldacena (2014); Bianchi et al. (2014, 2017, 2018b, 2018a). Solving this matrix model and deriving exact expressions for the observables that are valid for arbitrary values of parameters of the ABJM theory is an extremely difficult and challenging task.
@@
A powerful technique for addressing this problem is the Fermi gas approach Marino and Putrov (2012). In this framework, the ABJM matrix model is reinterpreted as a one-dimensional quantum gas of interacting fermions, with the Chern–Simons level $k$ playing the role of an effective Planck constant, $\hbar=2\pi k$. This approach has been successfully applied to the computation of the partition function $Z(N,k)$ of the ABJM theory on the round three-sphere $S^{3}$, see Hatsuda et al. (2015); Marino (2017) for reviews. Remarkably, the generating function $\Xi(z,k)=\sum_{N\geq 0}z^{N}Z(N,k)$, which has the meaning of the grand canonical partition function of the Fermi gas, admits a compact representation as a Fredholm determinant of a certain integral operator.  [^2] It is noteworthy that various observables in four-dimensional maximally supersymmetric Yang-Mills theory can likewise be expressed in terms of a Fredholm determinant of an integrable Bessel kernel (see Beccaria et al. (2022) and references therein). This representation is particularly powerful, as it enables the computation of these observables for arbitrary values of the ’t Hooft coupling Bajnok et al. (2024, 2025b, 2025a). In the strong-coupling regime, corresponding to the limit $\hbar=2\pi k\to 0$, the Fredholm determinant can be evaluated using a semiclassical expansion. When supplemented with nonperturbative (instanton) corrections, this analysis leads to the expression for the partition function in terms of Airy function.
@@
The Fermi gas formalism can also be applied to the computation of supersymmetric Wilson loops. In the ABJM theory, one can define two distinct supersymmetric Wilson loops, $W^{1/6}$ and $W^{1/2}$, which preserve four and twelve superconformal symmetries, respectively, out of the total $24$ of the theory Drukker et al. (2008); Chen and Wu (2010); Drukker and Trancanelli (2010); Drukker and others (2020). The $1/6$ BPS Wilson loop is defined as the vacuum expectation value
@@
Both Wilson loops (1) and (2) depend on a nonnegative integer $n$, which specifies the winding number. For $n=0$, they satisfy the normalization condition $W_{0}^{1/6}=1$ and similarly $W_{0}^{1/2}=0$. As a consequence of (2), it is sufficient to focus on the computation of the $1/6$ BPS Wilson loop (1).
The semiclassical expansion of the Wilson loops (1) and (2) within the Fermi gas approach was derived in Klemm et al. (2013). Neglecting nonperturbative (instanton) corrections, the resulting expressions for $W_{n}^{1/6}$ and $W_{n}^{1/2}$ take the form of ratios of Airy functions and their derivatives. In parallel development, the $1/6$ BPS Wilson loop was computed numerically with high precision for a wide range of parameters $k$ and $N$ in Okuyama (2016a). These numerical studies revealed that, for winding number $n\geq 2$, the perturbative part of the $1/6$ Wilson loop differs from the semiclassical expression obtained in Klemm et al. (2013). This discrepancy was unexpected, especially given that the semiclassical expression for the partition function $Z(N,k)$ agrees perfectly with numerical data Hatsuda et al. (2012). The origin of the mismatch remained unexplained.
@@
In the Fermi-gas framework, the Wilson loop (1) admits an exact representation in terms of the resolvent of the same operator that appears in the Fredholm determinant expression for the partition function $Z(N,k)$. The corresponding Fredholm determinant is known to satisfy a system of coupled nonlinear TBA-like equations. These equations were originally introduced independently in the study of $\mathcal{N}=2$ models in two dimensions Cecotti et al. (1992) and of self-avoiding 2d polymers Zamolodchikov (1994), and were subsequently proven in Tracy and Widom (1996b).
Building on the method developed by Tracy and Widom in Tracy and Widom (1996b), we derive a set of exact equations governing the function $W_{n}^{1/6}$. We show that these equations are sufficiently powerful to determine the Wilson loops for arbitrary values of the parameters, without relying on the semiclassical approximation.
@@
We verified that the obtained results for the Wilson loops are in perfect agreement with the expressions obtained in Okuyama (2016a) from numerical fitting. Combined with the explicit expressions for the free energy, they provide an efficient method for computing the Wilson loops for arbitrary values of the parameters of the ABJM theory.
@@ 2 Fermi gas approach
Supersymmetric localization allows us to express the partition function of the ABJM theory on the three-sphere and the $1/6$ BPS Wilson loop (1) in terms of the same function $W_{n}=W_{n}(N,k)$. This function is given by a finite-dimensional integral Kapustin et al. (2010)
@@
To illustrate this, it is sufficient to consider the Abelian case with $N=1$. In this case, the integral in (5) can be evaluated explicitly, yielding Okuyama (2016a)
@@
This expression is regular for $k$ obeying (7), but it develops a singularity at $k=2n$. The singularity originates from the divergence of the integral (5) at this value of $k$. As we demonstrate below, the same singular behavior persists for arbitrary $N$.
@@ 2.1 Generating functions
As follows from the first relation in (6) these two functions are not independent and are related to each other as
@@
In a similar manner, the generating function $\mathcal{W}_{n}(z,k)$ has the meaning of the expectation value of the holonomy $\mathop{\rm tr}\nolimits(U^{n})=\sum_{i=1}^{N}e^{n\mu_{i}}$ in a grand canonical ensemble.
@@
Applying (6) and replacing the functions $Z(N,k)$ and $W_{n}(N,k)$ with their integral representation (5), we obtain rather complicated expressions for the generating functions (2.1). As was shown in Marino and Putrov (2012), these expressions can be cast into a compact form by noticing that the function $Z(N,k)$ admits a determinant representation
@@
In a similar way, the generating function $\mathcal{W}_{n}(z,k)$ defined in (2.1) can be written in the operator form Klemm et al. (2013); Hatsuda et al. (2013a),
@@
where the operator $\bm{U}=e^{(\bm{x}+\bm{p})/k}$ provides a representation of the holonomy in the grand canonical ensemble. We verify that for $n=0$ the relations (15) and (17) are in agreement with (10). For $n\neq 0$, the operator in (17) can be simplified using the Baker-Campbell-Hausdorff formula as $\bm{U}^{n}=e^{n\bm{p}\over k}e^{n(\bm{x}+i\pi n)\over k}$ leading to the following representation Okuyama (2016a)
@@
As we demonstrate below, relations (15) and (18) provide a powerful tool for evaluating un-normalized Wilson loop $W_{n}(N,k)$ for arbitrary positive $N$ and $k$. For the first few values of $N$, this function can be found from small $z$ expansion of the generating functions (2.1). Expanding the right-hand side of (15) and (18) in powers of $z$, we get from (2.1)
@@ 2.2 Semiclassical approximation
The advantage of considering this regime is that, within the Fermi gas framework, the small-$k$ behavior of the generating functions $\Xi(z,k)$ and $\mathcal{W}_{n}(z,k)$ defined in (2.1) can be systematically derived from the semiclassical WKB expansion of the quantum Fermi gas. It is important to emphasize, however, that it is not obvious a priori whether this expansion remains valid for finite values of $k$, or, equivalently, whether the semiclassical expressions can be analytically continued from small $k$ to positive integer values without encountering singularities. This question has been investigated in Hatsuda et al. (2012); Okuyama (2016a) by comparing semiclassical predictions with numerical evaluations of (15) and (18). The analysis shows that, while the semiclassical expansion accurately describes the function $\Xi(z,k)$, it fails for $\mathcal{W}_{n}(z,k)$ when $n\geq 1$.
@@
By combining the semiclassical expansion with numerical analysis and imposing certain nontrivial consistency conditions, the grand canonical partition function $\Xi(z,k)$ defined in (2.1) and (15) was conjectured to have the following form at large $z=e^{\mu}$ and finite $k$, see Hatsuda et al. (2015) for a review,
@@
where the constant term $A(k)$ admits an explicit integral representation Marino and Putrov (2012); Hanada et al. (2012).
@@
Furthermore, the grand canonical partition function (21) simplifies significantly for $k=1,2,4,8$ and the corresponding function $\Xi(N,k)$ can be written in a closed form in terms of the Jacobi theta functions Codesido et al. (2015); Grassi et al. (2016); Okuyama (2016b).
@@
Previous studies have shown that, at large $\mu$ and finite $k$, the function $\mathsf{W}_{n}(\mu,k)$ admits the following general representation Klemm et al. (2013); Hatsuda et al. (2013a); Okuyama (2016a)
@@
The perturbative contribution (29) has been computed using two complementary approaches: through the semiclassical expansion of the Fermi gas ensemble Klemm et al. (2013) and via high-precision numerical evaluation of the winding Wilson loop Okuyama (2016a). The corresponding constant term was found to be different
@@
The imaginary and real parts of the nonperturbative contribution to the $1/6$ BPS Wilson loop (30) were studied numerically in Hatsuda et al. (2013a); Okuyama (2016a) for the first few winding numbers, n=1,2,3,4n=1,2,3,4444In virtue of (2), the function $\mathop{\rm Im}\nolimits\mathsf{W}_{n}^{\text{np}}(\mu,k)$ coincides with the nonperturbative part of the $1/2$ BPS Wilson loop. Because of the normalization factor in (28), the function $\mathop{\rm Re}\nolimits\mathsf{W}_{n}^{\text{np}}(\mu,k)$ corresponds to the imaginary part of the $1/6$ BPS Wilson loop in the conventions of Okuyama (2016a). By evaluating $\mathsf{W}_{n}^{\text{np}}(\mu,k)$ numerically for some positive integer values of $k$, the exact values of the instanton coefficients $w_{m,\ell}(\mu)$ were guessed for the lowest values of $m$ and $\ell$ through numerical fitting. Building on this analysis, a number of conjectures were formulated concerning the properties of the nonperturbative function $\mathsf{W}_{n}^{\text{np}}(\mu,k)$.
@@ 3 Tracy-Widom approach
This approach was originally developed by Tracy and Widom Tracy and Widom (1996b), who proved the conjecture formulated in Cecotti et al. (1992); Zamolodchikov (1994), according to which the Fredholm determinant of an integral operator of the same type as in (15) satisfies a TBA-like system of nonlinear equations. The integral operator in (15) involves the kernel $\rho(x,y)$ defined in (14). It represents a special case of the more general class of kernels analyzed in Tracy and Widom (1996b) [^5] Note that the kernel $\rho(x,y)$ is not symmetric, $\rho(x,y)\neq\rho(y,x)$. To bring it to the form considered in Tracy and Widom (1996b), it has to be symmetrized by a similarity transformation, which leaves the Fredholm determinant (15) invariant.
We begin by introducing an auxiliary function which plays a central role in our analysis  [^6] In the notation of Tracy and Widom (1996b), this function is given by the sum of functions $Q(\theta)$ and $P(\theta)$ for $\theta=x/k$.
@@
The rationale for introducing the auxiliary function (32) is that the generating function (18) can be expressed in terms of this function
@@
Replacing the $\psi$-function in (3) with its small $z$ expansion (3), we can systematically expand the generating function $\mathcal{W}_{n}(z,k)$ in powers of $z$ and reproduce the relation (19). To find $\mathcal{W}_{n}(z,k)$ for arbitrary $z$, we have to determine the function (32). This will be done in two steps. First, we formulate a set of nontrivial relations the function $\psi(x|z)$ has to satisfy and, then, present the solution.